# Towards Supporting Legal Argumentation with NLP: Is More Data Really All You Need?

### v1

@@ 1 Introduction
Law has been an attractive domain for AI in both symbolic knowledge representation and statistical NLP. Both strands share the common goal of supporting legal practice through enhancing legal research, document analysis, drafting, and decision making. A focal question distinguishing them remains whether, and how, the process of legal reasoning underlying all textual data shall be explicitly represented or left to opaque components, such as generative language models or neural classifiers.
@@
AI&Law as a field started started in the 1970s, when Buchanan and Headrick (1970) suggested that computer modeling of legal reasoning would be a promising area for research to better understand legal reasoning and argumentation. Many approaches have been proposed over the past three decades capturing several types of reasoning by means of symbolic representations. Some 50 years after the field’s beginnings, the legal profession is experiencing considerable disruption by NLP technology, most prominently large language models (LLMs) Katz et al. (2023b). In this paper, we provide a review of AI&Law work offering faithful modeling of legal reasoning but also requiring expensive legal expertise. We contrast this to modern, largely non-explainable, data-driven methods, which predict legal conclusions directly without engaging in any explicit legal reasoning.
@@
  • LLMs widen the knowledge acquisition bottleneck for domain model construction.
@@ 2 Legal Systems in a Nutshell
The main vision of AI&Law is that state and private actors in all aspects of the legal system can benefit from supporting software that seamlessly connects to the concepts and concerns they have been trained for and work with. Notably, legal reasoning not only happens in courts, but also in public administration and law enforcement (i.e., the executive branch), where law needs to be applied to specific situations (e.g., permits, taxes, police, public safety, etc.). Human accountability is paramount for the trust in the overall workings of a democratically governed society. Hence, this vision is one of AI supporting human decision makers and not replacing or unduly influencing them.
@@ 3 Knowledge-based Approaches
Case-based Approaches The adversarial nature of law naturally demands to represent arguments for both dispute sides. The precedent-focused nature of US common law was a suitable domain for the development of what became known as ‘legal case based reasoning’ systems. In the prominent TAXMAN system, McCarty (1976) modeled the majority’s and minority’s theories and arguments in the famous tax law case of Eisner v Macomber, 252 U.S. 189 (1920) Eisner v. Macomber , surveying the intricacies one must account for if resolved to capture all decision-relevant concerns in depth. The HYPO system Ashley (1991) modeled parts of US Trade Secrets Law by means of dimensions. These are typical fact patterns that favor different sides of the dispute, and can be used to analogize and distinguish cases argumentatively by means of set comparison. The focal concept here is a ‘three-ply argument’: A proponent cites the most-on-point precedent with the greatest factor overlap. The opponent distinguishes by pointing to a disfavorable factor in the precedent but not in the new case, or a favorable factor in the current case but not in the precedent, and cites a counterexample precedent. Finally, the proponent offers a rebuttal by distinguishing the counterexample. This was built upon in the CATO system Aleven (1997), which arranged ‘factors’ into a hierarchy, on the basis of which more sophisticated argumentation was possible (e.g., using hierarchy parent factors).
@@
Berman and Hafner 1991 observe that the support of a precedent decision for a case to be argued is linked to its respective procedural setting. They distinguish the pleading, pre-verdict, and verdict stage. A further difference exists between decisions on procedural matters and decisions on matters of fact and/or law. A decision in favour of the defendant party based on a procedural matter (e.g., lack of evidence) may not support the same decision in a new case which shares the factual features of the precedent but is to be decided on its merits. The question of decision context has received limited attention in subsequent works (e.g., Wyner and Bench-Capon 2009; Verheij 2016). Even in recent work on NLP-based legal judgment prediction, case outcomes are often greatly simplified, up to the point of an impoverished binary variable of whether a party won the case or not.
@@
Abstract Dialectical Frameworks (ADFs) Brewka et al. (2013) generalize the AAF representation to node-and-directed-relations form with a set of local acceptance conditions. This allows both attack and support influence, resulting in an abstract yet intuitive model for legal reasoning. For example, the ANGELIC method Al-Abdulkarim et al. 2016b uses ADFs for representing case law in an explainable inference model on the basis of a hierarchical factor representation. The maintainability of such a representation is discussed in Al-Abdulkarim et al. (2016a). Atkinson et al. (2019) extended to reasoning about factors with magnitude, thereby going beyond purely boolean proposition representations of cases.
@@ 4 Data-driven Approaches
Towards Modern Legal NLP: Recent years saw a resurging interest in case prediction through the use of data-driven methods learning from the large datasets now available from different jurisdictions, such as the ECtHR Chalkidis et al. (2019, 2022a); Aletras et al. (2016); Medvedeva et al. (2021) Chinese Criminal Courts Luo et al. (2017); Yue et al. (2021), US Supreme Court Katz et al. (2017); Kaufman et al. (2019), Indian Courts Malik et al. (2021), French court of Cassation Şulea et al. (2017b, a), Supreme Court of Switzerland Niklaus et al. (2021), Turkish Constitutional court Sert et al. (2021), UK courts Strickson and De La Iglesia (2020), German courts Waltl et al. (2017), Brazilian courts Lage-Freitas et al. (2022) and Philippine courts Virtucio et al. (2018).
@@
Earlier works employed bag-of-words features Aletras et al. (2016); Şulea et al. (2017a, b); Virtucio et al. (2018); Shaikh et al. (2020); Medvedeva et al. (2020). More recent approaches use deep learning techniques Zhong et al. (2018, 2020); Yang et al. (2019) involving convolutional or recurrent networks followed by adoption of pre-trained transformer models Chalkidis et al. (2019); Niklaus et al. (2021), including legal-domain specific pre-trained variants Zheng et al. (2021); Chalkidis et al. (2020, 2023); Douka et al. (2021); Masala et al. (2021); Xiao et al. (2021); Hwang et al. (2022); Niklaus et al. (2023). Classification tasks on legal text interrelate, and so other words have leveraged dependencies between tasks for improving models Tyss et al. (2023b); Yue et al. (2021); Valvoda et al. (2023); Zhong et al. (2018); Feng et al. (2022); Ma et al. (2021); Dong and Niu (2021); Yang et al. (2019); Huang et al. (2021); Hu et al. (2018) and added additional loss constraints (such as contrastive learning exploiting label information), Tyss et al. (2023b); Zhang et al. (2023); Gan et al. (2022); Liu et al. (2022) and injected legal knowledge Tyss et al. (2023a); Gan et al. (2021); Zhong et al. (2020); Feng et al. (2022)
@@
Overall, one can observe a trend towards applying NLP models to legal text with little to no architectural bias or explicit domain representation. These are then compared along quantitative metrics, typically with regard to high level classification/prediction goals (e.g., case outcome variables and document-level keywords) at the cost of interpretability. As Berman & Hafner have observed in the 1990s, however, case outcomes are highly contextual in time, procedure, and socio-legal purpose. Classification benchmarks risk decoupling a sense of technical progress towards a notion of model ‘understanding’ from supporting a realistic task (e.g., legal argumentation) by focusing on a highly reductive representation of its outcome.
@@
Shift to Generative Models: (LLMs have also been evaluated against case outcome classification as a benchmark. Chalkidis (2023); Vats et al. (2023); Trautmann et al. (2022); Shui et al. (2023) tested various early models and found them to score relatively low in quantitative metrics, which stands in contrast to their scores on some bar exams Katz et al. (2023a); Freitas and Gomes (2023). They report on experiments with several models and prompting techniques, including zero/few-shot prompting, prompt ensembling, chain-of-thought, and activation fine-tuning. Yu et al. (2022, 2023) employ prompts that are derived from legal reasoning methods (such as the common law IRAC (Issue, Rule, Application, Conclusion). Trautmann (2023) uses prompt chaining with an initial summarization step to deal with lengthy legal documents. Jiang and Yang (2023); Deng et al. (2023) develop syllogism prompting providing the three deductive reasoning steps for major premise (article/law retrieval), minor premise (element extraction from facts) and conclusion (judgement).
LegalBench Guha et al. 2023 recently presented the first aggregated benchmark beyond classification-like evaluation to test the reasoning abilities of generative models. Kang et al. (2023) applies the IRAC methodology comprehensively to LegalBench subtasks. While ancillary challenges remain (e.g., the need to manually assess model performance non certain tasks), this development is in line with our arguments in this paper.
@@ 5 Challenges & Future Directions
Domain Model Construction Rule-based models of the law are powerful tools to develop software that supports legal practice, but constructing them demands considerable legal expertise. Modern LLMs put us into a position to create these structures in a (semi-) automated fashion. Savelka et al. (2023) shows constructive evidence of this, but it remains an open questions whether LLMs can systematize large complexes of legal source material into well-formed, legally correct representations. Ascribing factors to facts from text in unseen cases by means of developing classifiers requires training data relative to an exhaustively defined list of factors. The more likely scenario is that generative models can be prompted with specific facts to subsume them under a factor pattern description. For example, Gray et al. (2024) applied generative AI automatically to identify factors in Fourth Amendment auto stop cases.
@@
Role of Evaluation: Legal databases are more than conveniently available repositories of large quantities of text. Legal texts should be seen as tools in performing use case tasks, including, for example, using prior decisions in constructing and responding to arguments. This implies that legal NLP’s efforts should be evaluated in terms of how well they support such use cases Ashley (2022). Comparatively few legal NLP papers specify a use case as an evaluative framework, rely on evaluations by legal experts, or benchmark their automatic metrics against human evaluations. This should change.
@@ 6 Conclusions
This paper focuses on legal NLP as applied to tasks that involve the application of legal source material to case facts, analysis of case texts, and legal argumentation in general. Other subfields of NLP in the legal domain do not focus on argumentation about the legal significance of case facts, such as technology-assisted review in e-Discovery, contract analysis, and patent search. Similarly, legal question answering, automatic summarization of judgments, legal information retrieval, and models supporting regulatory compliance, although important, are in focus for our argumentation-related narrative.
We strive to synthesize a very broad notion of the important role of expert legal knowledge to facilitate better NLP systems that will be of high utility to the stakeholders involved in the ecosystem. In our way, the way forward requires input from diverse perspectives and collaboration across multiple disciplines, including law, computer science, linguistics, and ethics to achieve a comprehensive understanding of the challenges and opportunities. We hope that the insights provided in this paper will stimulate an open discussion within the legal NLP community and beyond.

### v3

@@ 1 Introduction
Law has been an attractive domain for AI in both symbolic knowledge representation and statistical NLP. Both strands share the common goal of supporting legal practice through enhancing legal research, document analysis, drafting, and decision making. A focal question distinguishing them remains whether, and how, the process of legal reasoning  [^1] By ‘legal reasoning’, we refer to the wide range of activities involving interpreting, arguing, and applying legal principles to reach conclusions. Legal reasoning is not a single task but a collection of related tasks around the main theme of legal decision-making as the interrelation of more of less well-defined rules and societal values with the facts of a specific case towards an outcome. Given the limited space available, we use ‘legal reasoning’ as an umbrella term to cover the diverse contributions on this topic in the literature. underlying all textual data shall be explicitly represented or left to opaque components, such as generative language models or neural classifiers.
@@
AI & Law as a field started started in the 1970s, when Buchanan and Headrick (1970) suggested that computer modeling of legal reasoning would be a promising area for research to better understand legal reasoning and argumentation. Many approaches have been proposed over the past three decades capturing several types of reasoning by means of symbolic representations. Some 50 years after the field’s beginnings, the legal profession is experiencing considerable disruption by NLP technology, most prominently large language models (LLMs). In this paper, we provide a review of AI & Law work offering faithful modeling of legal reasoning but also requiring expensive legal expertise. We contrast this to modern, largely non-explainable, data-driven methods, which predict legal conclusions directly without engaging in any explicit legal reasoning.
@@
  • LLMs help alleviate knowledge acquisition bottleneck for domain model construction.
@@
While prior surveys of Katz et al. (2023b) and Zhong et al. (2020b) focus on cataloging various use cases, tasks, and NLP techniques in legal AI, our paper critically examines the historical integration of expert knowledge into legal systems and advocates for its revival and synthesis with data-driven methods. We emphasize the unique value of expert-informed knowledge in ensuring legal reasoning aligns with established principles, which is not the primary focus of the aforementioned reviews. In contrast to Mahari et al. (2023), which highlights the disconnect between the tasks that are pursued in legal NLP research and the actual needs of legal practitioners, our work emphasizes the critical importance of integrating expert-informed knowledge to avoid this gap. We also present directions for synthesizing expert knowledge with current technological advancements, thereby overcoming traditional bottlenecks in knowledge acquisition and enhancing the efficacy of structured argumentation models.
Most importantly, we contribute a comprehensive distillation of the conceptual ideas developed and researched by the AI & Law community prior to the recently surging interest in law as an application domain for mainstream NLP. In part, our motivation is to connect these communities. Much legal NLP work does not build on formal models of legal knowledge and reasoning but characterizes it mostly as precursor work to modern statistical methods. Our position is that this view does not do justice to the insights gained and legal authenticity captured in this body of research. Symbolic AI & Law has thought about how to incorporate legal expertise in models much deeper than most current NLP works, and hence the fields should merge and learn from one another. We strive to drive home the necessity of a paradigm shift in legal NLP, one that values and integrates the profound expertise of domain specialists with the capabilities of data-driven technologies.
@@ 2 Legal Systems in a Nutshell
It is worthwhile to acknowledge that legal systems are inherently human-centric, involving complex decision-making processes where persuasion, interpretation, and subjective judgment play critical roles. Legal decisions are not solely about determining which side should win as a matter of justice, but about who can present the most convincing argument within the framework of established laws, principles, and precedents. The main vision of AI & Law is that state and private actors in all aspects of the legal system can benefit from supporting software that seamlessly connects to the concepts and concerns they have been trained for and work with. Notably, legal reasoning not only happens in courts, but also in public administration and law enforcement (i.e., the executive branch), where law needs to be applied to specific situations (e.g., permits, taxes, public safety, etc.). Human accountability is paramount for the trust in the overall workings of a democratically governed society. Hence, this vision is one of AI supporting human decision makers and not replacing or unduly influencing it.
@@ 3 Knowledge-based Approaches
Case-based Approaches The adversarial nature of law naturally demands to represent arguments for both dispute sides. The precedent-focused nature of US common law was a suitable domain for the development of what became known as ‘legal case based reasoning’ systems. In the prominent TAXMAN system, McCarty (1976) modeled the majority’s and minority’s theories and arguments in the famous tax law case of Eisner v Macomber, 252 U.S. 189 (1920) Eisner v. Macomber , surveying the intricacies one must account for, if resolved to capture all decision-relevant concerns in depth. The HYPO system Ashley (1991) modeled parts of US Trade Secrets Law by means of dimensions. These are typical fact patterns that favor different sides of the dispute, and can be used to analogize and distinguish cases argumentatively by means of set comparison. The focal concept here is a ‘three-ply argument’: A proponent cites the most-on-point precedent with the greatest factor overlap. The opponent distinguishes by pointing to a disfavorable factor in the precedent but not in the new case, or a favorable factor in the current case but not in the precedent, and cites a counterexample precedent. Finally, the proponent offers a rebuttal by distinguishing the counterexample. This was built upon in the CATO system Aleven (1997), which arranged ‘factors’ into a hierarchy, on the basis of which more sophisticated argumentation was possible (e.g., using hierarchy parent factors).
@@
Berman and Hafner 1991 observe that the support of a precedent decision for a case to be argued is linked to its respective procedural setting. They distinguish the pleading, pre-verdict, and verdict stage. A further difference exists between decisions on procedural matters and decisions on matters of fact and/or law. A decision in favour of the defendant party based on a procedural matter (e.g., lack of evidence) may not support the same decision in a new case which shares the factual features of the precedent but is to be decided on its merits. The question of decision context has received limited attention in subsequent works (e.g., Wyner and Bench-Capon 2009; Verheij 2016). Even in the recent works on NLP-based legal judgment prediction, case outcomes are often greatly simplified, up to the point of an impoverished binary variable of whether a party won the case or not.
@@
Abstract Dialectical Frameworks (ADFs) Brewka et al. (2013) generalize the AAF representation to node-and-directed-relations form with a set of local acceptance conditions. This allows both attack and support influence, resulting in an abstract yet intuitive model for legal reasoning. For example, the ANGELIC method Al-Abdulkarim et al. (2016b) uses ADFs for representing case law in an explainable inference model on the basis of a hierarchical factor representation. The maintainability of such a representation is discussed in Al-Abdulkarim et al. (2016a). Atkinson et al. (2019) extended to reasoning about factors with magnitude, thereby going beyond purely boolean proposition representations of cases.
@@ 4 Data-driven Approaches
Towards Modern Legal NLP: Recent years saw a resurging interest in case prediction through the use of data-driven methods learning from the large datasets now available from different jurisdictions, such as the ECtHR Chalkidis et al. (2019, 2022a, 2021); Aletras et al. (2016); Medvedeva et al. (2021); SAYS (2020); Tyss et al. (2023b, a); Santosh et al. (2024c); Liu and Chen (2017); Medvedeva et al. (2020); SAYS (2020) Chinese Criminal Courts Luo et al. (2017); Yue et al. (2021); Zhong et al. (2020a, 2018); Yang et al. (2019), , US Supreme Court Katz et al. (2017); Kaufman et al. (2019), Indian Courts Malik et al. (2021); Shaikh et al. (2020) French court of Cassation Şulea et al. (2017b, a); Bertalan and Ruiz (2020) Supreme Court of Switzerland Niklaus et al. (2021), Turkish Constitutional court Sert et al. (2021), UK courts Strickson and De La Iglesia (2020), German courts Waltl et al. (2017), Brazilian courts Lage-Freitas et al. (2022) and Philippine courts Virtucio et al. (2018).
@@
Earlier works employed bag-of-words features Aletras et al. (2016); Şulea et al. (2017a, b); Virtucio et al. (2018); Shaikh et al. (2020); Medvedeva et al. (2020). More recent approaches use deep learning techniques Zhong et al. (2018, 2020a); Yang et al. (2019) involving convolutional or recurrent networks followed by adoption of pre-trained transformer models Chalkidis et al. (2019); Niklaus et al. (2021), including legal-domain specific pre-trained variants Zheng et al. (2021); Chalkidis et al. (2020, 2023); Douka et al. (2021); Masala et al. (2021); Xiao et al. (2021); Hwang et al. (2022); Niklaus et al. (2023). Classification tasks on legal text interrelate, and so other words have leveraged dependencies between tasks for improving models Santosh et al. (2023a); Yue et al. (2021); Valvoda et al. (2023); Zhong et al. (2018); Feng et al. (2022); Ma et al. (2021); Dong and Niu (2021); Yang et al. (2019); Huang et al. (2021); Hu et al. (2018) and added additional loss constraints (such as contrastive learning exploiting label information), Tyss et al. (2023b); Zhang et al. (2023); Gan et al. (2022); Liu et al. (2022) and injected legal knowledge Liu et al. (2023); Santosh et al. (2023b, 2024c); Gan et al. (2021); Zhong et al. (2020a); Feng et al. (2022)
@@
Overall, one can observe a trend towards applying NLP models to legal text with little to no architectural bias or explicit domain representation. These are then compared along quantitative metrics, typically with regard to high level classification/prediction goals (e.g., case outcome variables and document-level keywords) at the cost of interpretability. As Berman & Hafner have observed in the 1990s, however, case outcomes are highly contextual in time, procedure, and socio-legal purpose. Classification benchmarks risk decoupling a sense of technical progress towards a notion of model ‘understanding’ from supporting a realistic task (e.g., legal argumentation) by focusing on a highly reductive representation of its outcome. For instance, case outcome predictions are often treated as binary targets based on the majority opinion, even though judges on the same bench frequently have conflicting reasoning, leading to dissenting or concurrent opinions Xu et al. (2024). This reductive approach overlooks the nuanced legal argumentation underpinning each decision, focusing on a single outcome instead of capturing the depth of legal reasoning and debate.
@@
Shift to Generative Models: LLMs have also been evaluated against case outcome classification as a benchmark. Chalkidis (2023); Vats et al. (2023); Trautmann et al. (2022); Shui et al. (2023) tested various early models and found them to score relatively low in quantitative metrics, which stands in contrast to their scores on some bar exams Katz et al. (2023a); Freitas and Gomes (2023). They report on experiments with several models and prompting techniques, including zero/few-shot prompting, prompt ensembling, chain-of-thought, and activation fine-tuning. Yu et al. (2022, 2023) employ prompts that are derived from legal reasoning methods (such as the common law IRAC (Issue, Rule, Application, Conclusion). Trautmann (2023) uses prompt chaining with an initial summarization step to deal with lengthy legal documents. Jiang and Yang (2023); Deng et al. (2023) develop syllogism prompting providing the three deductive reasoning steps for major premise (article/law retrieval), minor premise (element extraction from facts) and conclusion (judgement).
LegalBench Guha et al. (2023) recently presented the first aggregated benchmark beyond classification-like evaluation to test the reasoning abilities of generative models. Kang et al. (2023) applies the IRAC methodology comprehensively to LegalBench subtasks. While ancillary challenges remain (e.g., the need to manually assess model performance non certain tasks), this development is in line with our arguments in this paper.
@@ 5 Challenges & Future Directions
Domain Model Construction Rule-based models of the law are powerful tools to develop software that supports legal practice, but constructing them demands considerable legal expertise. Modern LLMs put us into a position to create these structures in a (semi-) automated fashion. Savelka et al. (2023) shows constructive evidence of this, but it remains an open questions whether LLMs can systematize large complexes of legal source material into well-formed, legally correct representations. Ascribing factors from facts text in unseen cases by means of developing classifiers requires training data relative to an exhaustively defined list of factors. The more likely scenario is that generative models can be prompted with specific facts to subsume them under a factor pattern description. For example, Gray et al. (2024) applied generative AI automatically to identify factors in Fourth Amendment auto stop cases.
@@
Role of Evaluation: The true value in NLP for legal applications lies in producing, structuring, and assessing arguments about legal conclusions in an explainable way so that they may maximally support human experts. This human-centric nature of legal systems introduces a level of complexity that purely data-driven systems often struggle to capture when classifying variables from close-to-raw data. By the same token, LLMs may generate text that reads lawyer-like, but integrating them in processes of legal practice regularly involves interfacing them with symbolic data structures on both input and output ends, as well as maximizing consistency and correctness of generated text in ways that is defined by the legal concepts of the application context. This may include obfuscating cumbersome and error-prone model prompting behind traditional user interfaces composed of elements that map to symbols in the domain (e.g., types of contract clauses, factor-like aspects of cases, information elements of interest to draft process memoranda, etc.). The complexity of human legal decision-making highlights the inadequacy of current evaluation metrics. Legal NLP works should, ideally, tangibly indicate progress towards optimal argumentation support systems for legal practitioners, yet frequently convenient evaluations are prioritized over informative ones. This is, of course, due to the nuanced and often ill-defined characteristic of legal practice tasks. Still, legal databases are more than large repositories of text for autoregressive pre-training, but resources tend towards tackling these use cases, including, for example, using prior decisions in constructing and responding to arguments. Legal NLP’s efforts should be evaluated - and reviewed - in terms of how well models provide such functionality Ashley (2022). Many legal NLP works specify use cases, yet few account for them in their evaluative framework by conducting studies with legal experts, or benchmark their automatic metrics against human evaluations. Research on evaluation criteria that better capture the practical utility of legal NLP systems in real-world settings should be among our top priorities.
@@ 6 Conclusions
This paper focuses on legal NLP as applied to tasks that involve the application of legal source material to case facts, analysis of case texts, and legal argumentation in general. Other subfields of NLP in the legal domain do not focus on argumentation about the legal significance of case facts, such as technology-assisted review in e-Discovery, contract analysis, and patent search. Similarly, legal question answering, automatic summarization of judgments, legal information retrieval, and models supporting regulatory compliance, although important, are in focus for our argumentation-related narrative. We strive to synthesize a very broad notion of the important role of expert legal knowledge to facilitate better NLP systems that will be of high utility to the stakeholders involved in the ecosystem. In our way, the way forward requires input from diverse perspectives and collaboration across multiple disciplines, including law, computer science, linguistics, and ethics to achieve a comprehensive understanding of the challenges and opportunities. We hope that the insights provided in this paper will stimulate an open discussion within the legal NLP community and beyond.